Assume the​ risk-free rate is 7.2​% and the expected return on the market portfolio is 9.3​%. Use the capital asset pricing model​ (CAPM) to find the required return for each of the securities in the table​ here

   Beta
A         1.26
B         0.97
C        0.21
D         1.07
E         0.65

The required return for investment A is _______​%. ​(Round to one decimal​ place.)

The required return for investment B is ______% ​(Round to one decimal​ place.)

The required return for investment C is ______​%. (Round to one decimal​ place.)

The required return for investment D is _______% ​(Round to one decimal​ place.)

The required return for investment E is _______​%.​(Round to one decimal​ place.)
Answer #1

required return=risk-free rate +Beta*(market rate- risk-free rate)

Required return
A 7.2+(9.3-7.2)*1.26=9.8%(Approx).
B 7.2+(9.3-7.2)*0.97=9.2%(Approx).
C 7.2+(9.3-7.2)*0.21=7.6%(Approx).
D 7.2+(9.3-7.2)*1.07=9.4%(Approx).
E 7.2+(9.3-7.2)*0.65=8.6%(Approx).